OpenStudy (anonymous):

Can someone check my answers please??? A plane intersects a double-napped cone so that it contains the generating line. Which conic section is formed? (Points : 3) point line circle parabola **** i think the answer is line

OpenStudy (bohotness):

why do you think its line?

OpenStudy (bohotness):

YOU ARE CORRCET XD

OpenStudy (anonymous):

yay thank you heres the next one

OpenStudy (bohotness):

okay

OpenStudy (anonymous):

What is the length of the segment whose endpoints are A(4, -3) and B(10, 6)? (Points : 3) Option A: 3 square root 5 Option B: 3 square root 6 Option C: square root 93 Option D: squre root 117 I think its D

OpenStudy (anonymous):

That isnt the same question as the one I have

OpenStudy (bohotness):

click veiw more

OpenStudy (anonymous):

I did its not the same

OpenStudy (bohotness):

2): the vector between the point is (10 - 5, 6 - -1) = (5, 7) its magnitude is √(5^2 + 7^2) = √74

OpenStudy (anonymous):

But we arent using the same numbers its the same question type but not using the same numbers

OpenStudy (bohotness):

hold on lovve

OpenStudy (anonymous):

k

OpenStudy (bohotness):

i agree

OpenStudy (anonymous):

A bicycle is drawn on a grid such that the front tire is represented by the equation (x - 36)^2 + y^2 = 16. What is the tire's diameter? (Points : 3) 16 8 6 4 I think its B

OpenStudy (bohotness):

tandard form of circle: (x - h)^2 + (y - k)^2 = r^2, where (h, k) is the center and r is the radius The tire's radius is √16 = 4, so its diameter is 2(4) = 8.

OpenStudy (anonymous):

Thanks so much for everything

OpenStudy (bohotness):

XD yw

OpenStudy (bohotness):

XD yw